OpenStudy (anonymous):

Find the vertices of : \[x ^{4}+y ^{4}=1\] vertices is:point where derivative of kurvature is 0

OpenStudy (unklerhaukus):

graph it

OpenStudy (anonymous):

:). I need analitic solution...

OpenStudy (unklerhaukus):

yeah well you can work out the details once you have the solution,

OpenStudy (anonymous):

i am more intrested in a proces then in a solution itself

OpenStudy (unklerhaukus):

well it might be less intuative

OpenStudy (anonymous):

hmm what do you mean by curvature is zero? how do you measure curvature?

OpenStudy (anonymous):

f''(x) = 0?

OpenStudy (anonymous):

curvature is a mesure of change of the direction of unit tangent to arc length. It's one of geometric invariants of a curve

OpenStudy (anonymous):

f''(x) = 0? no

OpenStudy (unklerhaukus):

have you taken the derivatives o the formula

OpenStudy (anonymous):

if it helps anyone: the curvature function can be obtained by implicit differentiation

OpenStudy (anonymous):

thinking of x ^{4}+y ^{4}=1 as F(x,y) = x ^{4}+y ^{4}-1=0 and y(x)

OpenStudy (anonymous):

ohh it's calc iii or beyond me :/ ...sorry @dumbcow @satellite73 maybe they know how to do it.

OpenStudy (anonymous):

my aprouch was: F(x,y(x)) F'(1,y') F''(0,y'') so curvature function wouold be: k=y''/[(1+y'^2)^(3/2)] but not sure if this is ok...

OpenStudy (anonymous):

@Zarkon ?

OpenStudy (anonymous):

any sugestions @dumbcow ?

OpenStudy (dumbcow):

do you have to use curvature function (k) ? wouldn't the vertices be where the slope is 0 at (0,+-1)

OpenStudy (dumbcow):

btw you are correct with curvature equation http://en.wikipedia.org/wiki/Curvature#Curvature_of_a_graph

OpenStudy (anonymous):

not really. If you look at the graph of this function http://www.wolframalpha.com/input/?i=x+%5E%7B4%7D%2By+%5E%7B4%7D%3D1 the slope is not 0 at the vertices.

OpenStudy (anonymous):

ya but i am getting nowhere to find the actual values.....

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x+%5E4%2By+%5E4-1%3D0 this is the right one, sry

OpenStudy (dumbcow):

\[y' = -\frac{x^{3}}{y^{3}}\] \[y'' = -\frac{3x^{2}y^{4}+3x^{6}}{y^{7}}\] correct?

OpenStudy (anonymous):

1º is correct, checking the 2º...:)

OpenStudy (dumbcow):

oh it can be simplified if you factor out a 3x^2 and x^4+y^4 = 1 \[y''=-\frac{3x^{2}}{y^{7}}\]

OpenStudy (anonymous):

y'' is correct i think , just i am getting oposit sign.... maybe i am wrong

OpenStudy (anonymous):

i am getting \[y''=3x ^{2}/y ^{7}\]

OpenStudy (dumbcow):

when you plugged in y' did you include the neg sign? it is neg though, look at implicit derivatives on wolfram

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but, it's ok continue. I will check my steps later

OpenStudy (dumbcow):

ok, if we now look at curvature function after plugging everything in and simplifying i get \[k = -\frac{3x^{2}y^{2}}{(x^{6}+y^{6})^{3/2}}\]

OpenStudy (anonymous):

so the x derivative of this should be = 0

OpenStudy (dumbcow):

ok that derivative was a little messy, after setting =0 i was able to get to here \[2(x^{10}-y^{10}) = 7x^{4}y^{4}(y^{2}-x^{2})\] there should be 4 solutions right?

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=2%28x^10-y^10%29+%3D+7x^4y^4%28y^2-x^2%29 according to wolfram solution is y = +-x \[\pm x = \pm \sqrt[4]{1-x^{4}}\] \[x = \pm \sqrt[4]{\frac{1}{2}} \]

OpenStudy (anonymous):

thx a lot @dumbcow